Please help me solve this exponents question:
2^-3 / 2^4
(In words: Two raised to negative three divide by two raised to the power of four).
Kindly show the steps.
Thanking you in advance.

Answer by jim_thompson5910(35256)

2^-3 / 2^4 = 2^(-3 - 4) = 2^(-7) = 1/(2^7) = 1/128



So 2^-3 / 2^4 = 1/128